# MeSH Concept Relevance and Knowledge Evolution: A Data-driven Perspective

### v4

@@ Abstract
[inst2]organization=Université Paris-Saclay, CNRS, Laboratoire Interdisciplinaire des Sciences du Numérique,city=Orsay, postcode=91400, country=France
1 Introduction
The vast amount and complexity of the scientific biomedical literature require sophisticated methods of knowledge structuring and reasoning [1, 2]. Knowledge organization systems (KOSs), such as terminologies, thesauri, and ontologies, have been designed and used to encode concepts to improve discoverability, knowledge representation, and reasoning in biomedical corpora [1]. Among those KOSs, the Medical Subject Headings (MeSH) vocabulary indexes, structures, and enables search in the most prominent scientific biomedical corpus, i.e., the MEDLINE database accessible via the PubMed interface [3].
MeSH is a controlled vocabulary maintained by the National Library of Medicine [4]. In essence, it is a thesaurus that includes different types of concepts [1, 4, 5]. Descriptors are the primary elements of indexing and retrieval and are used to annotate the MEDLINE/PubMed dataset. Qualifiers and supplementary records are also available to provide additional information concerning descriptors. MeSH headings denote concepts from the biomedical domain and serve as a standardized interpretation for a group of medical terms [1, 4]. These headings are hierarchically organized into 16 main categories and can be associated with corresponding descriptors, which allows a direct mapping of descriptors to their appropriate MeSH headings. In addition to supporting search and question-answering in the biomedical literature [6, 7], they help explore patterns characterizing the structure and evolution of the biomedical field, assist in evaluating research grants and awards, and identify research topics of high interest [8, 9, 10].
As the knowledge in the biomedical domain evolves, MeSH needs to be periodically modified and updated to reflect these changes [11, 12]. Based on the current knowledge state, MeSH concepts (i.e., headings) encode a certain amount of information, which might vary over time [13]. The relevance of some concepts may decrease or increase, making them obsolete or requiring them to change drastically or to be added [5, 13, 14]. As a consequence, the hierarchy structure can change. Thus, it is crucial to identify the most representative MeSH concepts [1, 15] to summarize the content of the thesaurus, prioritize actions for maintaining tasks in KOS concepts, and enhance information retrieval systems relying on KOS, such as PubMed [14, 16, 17].
Different approaches have been used to quantify the amount of information encoded in MeSH by analyzing various features, such as the hierarchy, bibliometric indices of articles, and the co-occurrence of descriptors [9, 5, 13, 18, 19]. The hierarchy was used to examine properties, and MEDLINE articles citing descriptors were analyzed using scientometric methods [13, 18]. Balogh et al. discovered that nodes in this thesaurus with many children are more likely to gain even more children [13]. Adding new concepts to the hierarchy creates a non-uniform distribution among the primary branches, and adding links follows a non-trivial preference. Chien et al. analyzed PubMed articles in medicine and health using the x-index [18]. Descriptors were used to analyze trends in general and personalized medicine [9]. The co-occurrence of descriptors and new descriptors was also investigated [5, 19]. Kastrin and Hristovski built a network using co-occurrences of descriptors to investigate indexing trends [19]. Nentidis et al. demonstrated that about a quarter of new descriptors in this thesaurus can be considered emerging concepts [5]. In contrast, the remainder of the concepts were derived from descriptors in previous releases. In particular, Ilgisonis et al. focused on related descriptors rather than purely annotated descriptors, i.e., descriptors that co-occurred with the annotated descriptors [9]. Despite some limited attempts to use instances, such as the occurrence of descriptors in PubMed articles, instance analysis remains largely unexplored for the quantification of MeSH concepts over time.
To bridge this gap, this study aims to investigate an approach to measure the relevance encoded by MeSH considering the instantiation of its concepts. To do so, we analyze three main factors - the usage of concepts in PubMed articles, the hierarchical structure of MeSH, and the citation network of PubMed articles - in view of information theory and network analyses. The approach includes four aspects to quantify the relevance of MeSH concepts: informativeness, usefulness, disruptiveness, and influence. To quantify the amount of information a concept conveys, i.e., its informativeness, we use entropy, the standard concept of information theory [20, 21]. We use category utility supported by the MeSH hierarchical structure to measure the usefulness of a concept [22]. Some concepts may be more disruptive due to different factors, defining their disruptiveness level. For instance, they can be used to annotate highly cited but consolidated technologies, not bringing novelty to the field [10, 23, 24]. Lastly, to quantify the influence of MeSH concepts, we use the centrality metric from graph theory [25], which can be used to quantify node importance in networks [26]. We applied these methods to the MEDLINE database, analyzing their MeSH annotations over time. We generated monthly data by mapping references to MeSH concepts and computing each metric for each concept. This data was then propagated to the MeSH hierarchy. Finally, results were combined using a ranking fusion method - reciprocal ranking fusion (RRF) - to aggregate the multidimensional scores and provide a relative relevance of the MeSH concepts.
The paper presents three main contributions. First, we propose an algorithm to compute the relevance of a MeSH concept by propagating information from a network of data instances to the KOS structure. The network of MeSH instances is based on the PubMed citation network, while the knowledge structure relies on the thesaurus hierarchy. Second, we assess the relevance of MeSH concepts quantified by multidimensional metrics - informativeness, usefulness, disruptiveness, and influence - and show how they vary at different hierarchical levels over time. Lastly, we provide a quantitative evaluation of the proposed approach by analyzing changes in the MeSH terminology and retracted papers.
2 Materials and Methods
We explore the evolution of MeSH by analyzing concept instantiations over different years in PubMed. For this purpose, we use PubMed/MEDLINE corpus, which contains articles manually annotated with MeSH descriptors. In this section, we explain the methodology used to create the dataset[dataset] [27]. We describe the metrics used to quantify relevance in a KOS. Then, we discuss strategies for propagating metrics to the hierarchy and computation of a metrics fusion. To conclude this section, we detail the evaluation criteria used in this work.
2.1 Dataset and pre-processing
We conducted our study using the 2022 collection of PubMed [^1] https://lhncbc.nlm.nih.gov/ii/information/MBR/Baselines/2022.html, which includes 33,403,054 articles collected until December 2021 (see the distribution of PubMed articles in Figure 7). PubMed indexes new biomedical and life science articles whenever they are published. Domain experts annotate articles in PubMed using MeSH descriptors, which are organized hierarchically into 16 categories represented by single letters, such as A for Anatomy, B for Organisms, and C for Diseases. The MeSH hierarchy’s complete list of categories and their respective numbers of instances per MeSH version are presented in Table 1.
Table 1: Number of articles (in millions) that map to different categories in the MeSH hierarchy over time.
Code Category label 2014 2015 2016 2017 2018 2019 2020 2021
A Anatomy 9.6 10.0 10.4 10.7 11.1 11.5 11.9 12.3
B Organisms 19.7 20.7 21.6 22.5 23.5 24.4 25.4 26.3
C Diseases 11.6 12.2 12.7 13.3 13.9 14.5 15.1 15.7
D Chemicals and Drugs 12.8 13.3 13.9 14.4 15.0 15.5 16.1 16.6
E
Analytical, Diagnostic and
Therapeutic Techniques, and
Equipment
15.4 16.1 16.9 17.6 18.3 19.1 19.8 20.5
F Psychiatry and Psychology 3.5 3.8 4.0 4.2 4.4 4.6 4.8 5.0
G Phenomena and Processes 12.9 13.5 14.1 14.7 15.3 15.9 16.6 17.1
H Disciplines and Occupations 2.8 2.9 3.0 3.2 3.3 3.4 3.6 3.7
I
Anthropology, Education,
Sociology, and Social
Phenomena
2.2 2.3 2.5 2.6 2.7 2.9 3.0 3.2
J
Technology, Industry, and
Agriculture
1.9 2.0 2.2 2.3 2.4 2.6 2.7 2.8
K Humanities 0.5 0.6 0.6 0.6 0.6 0.7 0.7 0.7
L Information Science 2.5 2.7 2.8 3.0 3.1 3.2 3.4 3.5
M Named Groups 7.7 8.1 8.5 8.9 9.4 9.8 10.2 10.6
N Health Care 8.3 8.8 9.4 9.9 10.4 11.0 11.6 12.1
Z Geographicals 2.7 2.9 3.1 3.3 3.5 3.7 3.9 4.1
For this collection, we built citation networks using PubMed article citations and the [dataset]iCite [^2] https://nih.figshare.com/articles/dataset/iCite_Database_Snapshot_2022-04/19763386 dataset [28] for a more comprehensive source of citations. We then created accumulative monthly networks, i.e., including nodes from previous months, from January 2014 until December 2021. To build the corresponding MeSH instance hierarchy, we used the MeSH descriptors assigned to articles in a particular month mapped to their respective MeSH tree code (See Figure 1d for an example). The first monthly citation network (January 2014) has 18,728,651 nodes and 381,296,042 edges, and as citation networks accumulate articles over time, the last one (December 2021) has 27,728,971 nodes and 673,415,629 edges. Due to the prohibitive time required to compute certain metrics, we used 10% of nodes randomly sampled from the original monthly citation network, and edges were included only when both nodes existed in the sampled network. This resulted in the first sampled citation network (January 2014) having 1,872,866 nodes and 3,807,856 edges, while the last sampled network (December 2021) had 2,772,898 nodes and 6,745,896 edges.
2.2 MeSH relevance quantification
Analogous to the relevance definition in information science and information retrieval [29], we define the relevance of a concept as how well a concept or set of concepts meets the representation needs of a dataset [30]. Similar to the definition of relevance provided by Saracevic [31], in our study, we consider multiple aspects to quantify the relevance of MeSH concepts, namely informativeness, usefulness, disruptiveness, and influence (Figure 1a) [^3] https://zenodo.org/records/13635979. Each article in MeSH is annotated with MeSH descriptors. Furthermore, descriptors map to one or more tree nodes (Figure 1c and Figure 1d). In the following, we describe these metrics and how the relevance of a MeSH concept is quantified according to them.
Figure 1: Aspects to quantify the relevance of MeSH concepts. a) Metrics quantifying each aspect. b) Intuition of computation of each metric in this work. c) Flow of mappings over time: PubMed articles are annotated with MeSH descriptors, which link to one or more nodes in the MeSH hierarchy. d) Example of mapping between MeSH descriptors and the MeSH hierarchy of the article “Approaches towards fighting the COVID-19 pandemic (Review).” Bottom: MeSH descriptors. Top: MeSH hierarchy of the example’s article, showing only relevant nodes for this case.
2.2.1 Informativeness
To quantify the relevance of a concept in terms of informativeness, we use the entropy measure introduced by Shannon [20, 21]. Entropy quantifies the level of uncertainty of a variable. In our case, if a concept is used frequently to annotate multiple papers, it will not be informative, as many other articles also contain this annotation. For example, the concept “Humans” (D006801) is used to annotate approximately 70% of PubMed articles. Therefore, this concept adds little information when it appears in an article, having a low entropy value. On the other hand, the concept “Proteins” (D011506) is used to annotate only about 2% of PubMed articles, which makes it much more informative.
As shown in Equation 1, we follow Shannon’s entropy definition $H(X)$, where $p(x)$ represents the probability of an event’s occurrence and $x$ refers to all possible values that a variable $X$ can take:
$H(X)=-\sum_{x}p(x)\log_{2}p(x).$ (1)
In our study, an event is represented by the instantiation of a MeSH concept, that is, by articles being mapped to nodes in the MeSH hierarchy, representing the concept’s usage (dotted red arrows in Figure 1). The probability of an event $p(x)$ is then calculated as the probability that a concept $x$ is mapped within all mappings in its hierarchical level. Mappings in the lower levels are propagated to their parents when calculating the latter’s probability.
2.2.2 Usefulness
To quantify relevance in terms of usefulness, we use the category utility measure to identify basic-level categories in a taxonomy [22]. Basic-level categories are preferred categories in natural and artificial concept hierarchies, and they carry most information about the category’s attributes [22, 32] (e.g., when shown an image of a particular oncological exam, most professionals will identify it as a cancer rather than as a disease or a colon cancer). A node in a hierarchy is useful to the extent that it can be expected to improve the ability of a person to accurately predict the features of instances of that category. As shown in Figure 1b, in this study, the mapping of articles to the hierarchy characterizes concepts. Therefore, a concept that belongs to a highly mapped branch in the hierarchy will be more useful (e.g., as node c15 in the example).
Category utility combines the frequency of a category’s occurrence with the probability of its features [22]. Let us define a category $c$ as a finite set of instances: $c=\{o_{1},o_{2},...,o_{n}\}$, where each instance $o_{i}$ can be described by a finite set of discrete features $F=\{f_{1},f_{2},...,f_{m}\}$. The category utility (CU) of category $c$ with a set of features $F$ can then be calculated as in Equation 2, where $p(c)$ is the probability of a category (i.e., of a MeSH concept in our case), $p(f_{k})$ is the probability of feature $f_{k}$ and $p(f_{k}|c)$ is the conditional probability of feature $f_{k}$ given category $c$:
$CU(c,F)=p(c)\sum_{k=1}^{m}[p(f_{k}|c)^{2}-p(f_{k})^{2}].$ (2)
To calculate category utility for the nodes in the MeSH hierarchy, we use mapping to articles $a_{i}$ as the set of features $F=\{a_{1},a_{2},...,a_{m}\}$, where each feature $a_{i}$ has a binary value (annotated or not with a category). To compute the probabilities, we then created a binary matrix $M_{ij}$ of $n$ tree nodes and $m$ features as follows:
$M_{ij}=\begin{cases}1,&\text{tree node $i$ is mapped by article $j$}\\ 0,&\text{otherwise}\end{cases}.$
2.2.3 Disruptiveness
To measure the relevance of a concept in terms of the disruptiveness aspect, we use the disruption index [23]. This index was originally designed to quantify the extent to which an invention (a concept in our case) consolidates or destabilizes the subsequent use of the components (referenced concepts in our case) on which it builds [23]. To compute the disruption index $D$ for a focal node, we can consider the citation network as a tripartite graph $G=(V_{1},V_{2},V_{3},E)$, where the set of nodes ${V_{1},V_{2},V_{3}}$ represents the articles, which are connected by directed edges $E$, representing their citations [25, 26]. Given a focal node $a$ (light red in Figure 1b), corresponding to an article for which we want to compute the level of disruptiveness, there are $j$ articles that reference article $a$ and at least one of its predecessors (purple box), $i$ articles reference $a$ but none of its predecessors (green box), and $k$ articles do not reference $a$ but at least one of its predecessors (dark gray box), disruption $D$ for $a$ can then be defined by Equation 3:
$D=\frac{n_{i}-n_{j}}{n_{i}+n_{j}+n_{k}},$ (3)
where $n_{i}$, $n_{j}$, and $n_{k}$ represent the number of $i$, $j$, and $k$ articles, respectively. The disruption score $D$ of an article is then propagated to the concepts mapped to it ($c2$, $c3$, and $c4$ in the example). Finally, the disruption index $D_{c}$ for a concept $c$ is the sum over the disruption scores for all articles $a_{1}$, , $a_{n}$ which are mapped to $c$ divided by the number of articles in the citation network, i.e., $D_{c}=\sum_{i=1}^{n}D_{i}/m$, where $D_{i}$ is the disruption index of the $i$th article mapped to $c$ and $m$ is the number of articles in the citation network. To compute the $D$ index, we used Figueiredo and Andrade’s disruption implementation [23].
2.2.4 Influence
In graph theory, the influence of a node is quantified through graph centrality, which is based on the importance of its neighboring nodes. We use this concept to measure the influence aspect of a MeSH concept. Instead of considering the MeSH hierarchy, which captures only ontological relations, we use the instances (i.e., mapped articles) to identify the influence of MeSH concepts. Similar to the disruptiveness aspect, we consider the citation network to compute centrality. In Figure 1b, we show a small citation network example to demonstrate how nodes (articles) influence their neighbors. An article, annotated with one or more concepts, denoted by $c_{1}$, ..., $c_{n}$, gains more influence based on the influence of its incoming articles (e.g., light red box).
In our work, we utilized the PageRank algorithm to compute the centrality of a node [33, 25], which uses the importance of the node’s neighbors to compute the node’s importance. The mathematical definition of PageRank of node $i$ in a directed graph $G$ is $x_{i}$ provided by Equation 4:
$x_{i}=\alpha\sum_{j}A_{ij}\frac{x_{j}}{k_{j}^{out}}+\beta,$ (4)
where the parameter $\alpha$ regulates the contribution between the eigenvector and the constant term $\beta$. The term $\beta$ is used to initialize nodes with zero out-degree. The adjacency matrix of $G$ is denoted by $A_{ij}$, which captures the edges between pairs of nodes (citations in our case). The out-degree of node $j$ is represented by $k_{j}^{out}$. Similarly to disruptiveness, the centrality $x_{i}$ of an article $i$ is propagated to the concepts it is mapped to, and the overall centrality $x_{c}$ of a code is computed by summing over all the centralities of articles that concept $c$ is mapped to divided by the number of articles in the citation network, i.e., $x_{c}=\sum_{i=1}^{n}x_{i}/m$, where $x_{i}$ is the centrality score of the $i$th article mapped to $c$ and $m$ is the number of articles in the citation network. We used the implementation PageRank algorithm from the NetworKit [^4] https://networkit.github.io/ library to calculate the graph centrality, where $\alpha=0.85$ and $\beta=1-\alpha$.
2.3 Propagation of data instances to the knowledge representation
For the different concept relevance aspects, we encoded the usage of MeSH instances and propagated them to the knowledge representation. Given the “is a” relationship between MeSH concepts, we have considered two methods for propagating information to higher-level nodes: direct mapping and a cumulative weighted propagation algorithm. For the information-theoretical measures (entropy and category utility), which are computationally efficient, we propagate the mapping of instances directly to their parent concepts and compute the relevance scores for each node in the hierarchy. For example, if a paper is annotated with the MeSH concept “Heart Diseases [C14.280]”, its parents “Cardiovascular Diseases [C14]” and “Diseases [C]” are also considered to be mapped to that paper. Then, the information-theoretical metrics are computed normally for the parent nodes (e.g., C14 and C), including their direct mappings as well as the propagated mappings based on the “is a” relationship.
We approach the network analysis metrics (disruption and graph centrality) differently due to their high computational complexity. Instead of combining the mappings into higher-level nodes, we aggregate scores of lower-level nodes into the scores of their parents. The assumption is that if a node is disruptive or influential, so will its parent. As described in Algorithm 1, the propagation method requires the MeSH hierarchy $H$ and the network-based relevance scores computed using the citation network (variable $tnm$) as input. In line 2, we obtain tree nodes per level of specificity using the hierarchy $H$; for instance, tree nodes C and D are in level 1, while tree nodes C01 and D01 are in level 2. Then, we iterate over the hierarchy from bottom to top, using a reversed level list (variable $llr$, line 3) to propagate the relevance scores of the lower-level concepts to the higher levels. First, at each level, we get a set of unique parents (lines 5-9). Then, we retrieve their children (line 11) and, utilizing their initial measurements (lines 14-18), compute the weighted parent relevance, variable $gmp$, based on the children’s relevance and the number of nodes in their level (lines 19 and 21). Finally, the parent value is calculated taking into account the direct mapping to its node in addition to the weighted children measurements (lines 22 and 23).
Algorithm 1 Propagation of graph-based metrics to the MeSH hierarchy
0:  $H$ (hierarchy), $tnm$ (graph metric values initialized from descriptors)
0:  $gmh$ (graph metric of tree nodes in the hierarchy)
1:  $gmh\leftarrow\{\}$
2:  $dpt\leftarrow get\_nodes\_in\_levels\_specificity(H)$
3:  $llr\leftarrow\text{sorted list of keys in }dpt\text{ in reverse order}$
4:  for each $current\_level$ in $llr$ do
@@ 2.3 Propagation of data instances to the knowledge representation
6:     for each $treenode$ in $dpt[current\_level]$ do
@@
8:        add $parent$ to $unique\_parents$
@@
10:     for each $parent$ in $unique\_parents$ do
@@
13:        for each $child$ in $children$ do
@@
15:              if $child$ has no children then
@@ 2.4 Consolidating the relevance ranking of MeSH codes
Each of the metrics described previously measures different relevance aspects of a MeSH concept according to its data instances. To have a unified relevance score, we propose to fuse the multidimensional aspects into a single rank. Inspired by ranking fusion methods in the information retrieval field, we consider that each concept $d\in D$, where $D$ is the set of MeSH codes, is ranked according to a relevance dimension $r\in R$, where $R$ is the set of four aspects considered in this work: informativeness, usefulness, disruptiveness, and influence.
The field of information retrieval proposes various techniques for merging rankings, such as CombSUM, CombMNZ, CombANZ, rank-biased precision, and reciprocal rank fusion (RRF) [34]. Due to its simplicity, effectiveness, and robustness  [34, 35], we use RRF for computing the overall relevance of a MeSH concept based on their instantiations. RRF is an unsupervised approach that combines the outcomes of different information retrieval methods with varying relevance scores into a single result set using the rank information. The mathematical definition of RRF is provided in Equation 5:
$RRF\_score(d\in D)=\sum_{r\in R}\frac{1}{k+r(d)},$ (5)
where, in our case, a concept $d$ belongs to the set of MeSH codes $D$, and $r(d)$ represents the rank of a code $d$ according to a metric $r$ from the set of metrics $R$. The constant $k$ minimizes the impact of outliers.
@@ 2.5 Evaluation criteria
We qualitatively analyse the trends of MeSH concepts according to the different relevance aspects for various levels of the hierarchy and versions from 2014 to 2021. To quantitatively evaluate the effectiveness of our approach, we conducted two experiments. First, we assess the relevance of concepts that changed compared to those that remained unchanged between MeSH versions. Second, we compared the relevance of concepts between retracted and non-retracted papers in PubMed. We used the Mann-Whitney test to compare relevance changes; p-values $<$ $0.05$ are considered statistically significant.
@@ 3 Results
In this section, we provide the results of assessing the relevance of MeSH concepts according to the multi-aspect model previously described. We start with the analysis results of the MeSH terminology evolution, highlighting the most relevant concepts at the different levels of the hierarchy. Then, we provide quantitative analyses of concept relevance in two use cases: i) terminology evolution and ii) paper retraction. For the first, we compare the relevance of concepts that changed from one version to another versus those that remained unchanged. For the latter we compare the relevance of concepts used to annotate retracted papers versus those used to annotate non-retracted papers.
@@ 3.1 MeSH concept relevance trends
We calculated the relevance of MeSH concepts from 2014 to 2021 using the PubMed 2022 snapshot annotated with the MeSH 2022 version. Relevance scores for the different aspects - disruptiveness, influence, informativeness and usefulness - as well as for the overall concept relevance were computed monthly. Figure 2a depicts the cumulative trends (i.e., aggregate) for the first level of the MeSH hierarchy (A, …, Z). Disruptiveness, influence, and usefulness decrease over time ($p-value<0.001$). Park and colleagues also corroborate the decreasing trend of disruptiveness, although their analysis was based on scientific articles, in contrast to ours, which is based on MeSH concepts [36]. On the other hand, informativeness, which measures the degree of uncertainty, tends to increase over time ($p-value<0.001$). Therefore, the degree of uncertainty also increases as more recent knowledge is added to PubMed.
@@
Figure 2: Trend of metrics relevance of MeSH concepts at level 1 and correlation of descriptors. a) The sum of MeSH categories per relevance aspect over time. b) MeSH categories per relevance aspect over time. c) Ranking of concepts at level 1 over time. d) Correlation of aspects. Column names are dis: disruptiveness (disruption), infl: influence (centrality), info: informativeness (entropy), use: usefulness (category utility), rel: concept relevance (RRF fusion). e) Trend of top concepts in categories C, D, and E at level 2 per metric.
@@
Figure 2b shows how the first level of concepts in MeSH varies over time for each relevance aspect. All of the categories show a decreasing trend ($p-value<0.001$). The most useful concepts (Figure 2b - bottom right) in the first level are D (Chemicals and Drugs), followed by N (Health Care) and C (Diseases), indicating they encode the basic level categories of PubMed. In influence, all categories show a decreasing trend ($p-value<0.001$). Concept C is the most influential (Figure 2b - top right), indicating that ‘diseases’ is a central concept in PubMed. It is followed by the concepts D and G (Phenomena and Processes), which include biological, chemical, and physiological phenomena descriptors.
@@
There is a group of non-medical concepts at the bottom in the trend of influence, grouping categories I (Anthropology, Education, Sociology, and Social Phenomena), N, H (Disciplines and Occupations), K (Humanities), M (Named Groups), L (Information Science), J (Technology, Industry, and Agriculture), Z (Geographicals) which present a more linear trend near 0. Trends of categories in disruption are mainly decreasing ($p-value<0.001$), though categories M and H are not statistically significant. J shows an increasing trend ($p-value<0.001$). On the other hand, K presents a growing trend, and I present no trend with no statistically significant $p-value$. Concept C is also the most disruptive in the first level of the MeSH hierarchy (Figure 2b - top left), being more than two times more disruptive than the top-2 concept (D). On the other hand, non-medical categories I, N, H, K, M, L, J, and Z are located at the bottom of this figure in an almost linear trend near 0, similar to the influence aspect. Trends of entropy are primarily increasing with a $p-value<0.001$. However, categories H, G, B, A, and D present a decreasing trend ($p-value<0.001$). Category C shows no trend (not statistically significant). Concept D has the highest informativeness score (Figure 2b - bottom left) followed by concept E (Analytical, Diagnostic and Therapeutic Techniques, and Equipment). However, it gradually decreases, with a standard deviation of 0.003377. The trends of this aspect remain constant over time, but this is due to the range of values of each MeSH category.
@@
Figure 2c shows the ranking of concepts in level 1. Most of the concepts retain their ranking over time, except for Diseases (C), Health Care (N), Disciplines and Occupations (H), and I, which show some variation. Chemicals and Drugs (D) consistently ranks higher than other MeSH categories. The second position in this ranking alternates between ‘Diseases’ and ‘Health Care’, with the latter prevailing over the former for five out of eight years.
@@
Figure 2e shows the top concept trend in C, D, and E categories per metric, i.e., at level 2 in the hierarchy. Most of the trends are decreasing over time. For disruptiveness, concepts C04 (Neoplasms), D10 (Lipids), and E04 (Surgical Procedures) decrease over time ($p-value<0.001$). In influence, concepts C04, D05 (Macromolecular Substances), and E04 have a decreasing trend with a $p-value<0.001$. Concerning informativeness, C23 (Pathological Conditions, Signs and Symptoms) and E05 (Investigative Techniques) show an increasing trend, while D12 (Amino Acids, Peptides, and Proteins) has a decreasing trend ($p-value<0.001$). In usefulness, concepts C12 (Urogenital Diseases), D12, and E06 (Dentistry) present decreasing trends with a $p-value<0.001$. The increasing trend of concept C23 may mean that the manifestation of diseases is changing over time, producing uncertainty. On the other hand, the decreasing informativeness trend of concept D12 can mean that knowledge in this tree of MeSH is more established as this concept is used more frequently, becoming less informative. The trend of concept E05 shows an increasing degree of uncertainty until 2019 when the uncertainty decreases. The full description of the concepts, i.e., MeSH headings, can be found in Table 3

### v5

@@ 2.3 Propagation of data instances to the knowledge representation
[…]
4:for each current_levelcurrent\_level in llrllr do
@@
6:  for each treenodetreenode in dpt[current_level]dpt[current\_level] do
@@
8:   add parentparent to unique_parentsunique\_parents
@@
10:  for each parentparent in unique_parentsunique\_parents do
@@
13:   for each childchild in childrenchildren do
@@
15:     if childchild has no children then
@@ 2.4 Consolidating the relevance ranking of MeSH codes
Each of the metrics described previously measures different relevance aspects of a MeSH concept according to its data instances. To have a unified relevance score, we propose fusing the multidimensional aspects into a single ranking. Inspired by ranking fusion methods in information retrieval, we consider that each concept $d\in D$, where $D$ is the set of MeSH codes, is ranked according to a relevance dimension $r\in R$, where $R$ is the set of four aspects considered in this work: informativeness, usefulness, disruptiveness, and influence.
Information retrieval research proposes various techniques for merging rankings, such as CombSUM, CombMNZ, CombANZ, rank-biased precision, and reciprocal rank fusion (RRF) [10]. Due to its simplicity, effectiveness, and robustness  [10, 41], we use RRF to compute the overall relevance of a MeSH concept based on its instantiations. RRF combines the outcomes of different information retrieval methods with varying relevance scores into a single result set using the rank information. The mathematical definition of RRF is provided in Equation 5, where, in our case, a concept $d$ belongs to the set of MeSH codes $D$, and $r(d)$ represents the rank of a code $d$ according to a metric $r$ from the set of metrics $R$. The constant $k$ minimizes the impact of outliers. We set $k$ to $60$, as defined in the original RRF paper[10, 41].
$\text{RRF}(d\in D)=\sum_{r\in R}\frac{1}{k+r(d)},$ (5)
@@ 2.5 Evaluation criteria
We qualitatively analyze the trends of MeSH concepts according to the different relevance aspects for various levels of the hierarchy and across MeSH versions from 2014 to 2021. To quantitatively evaluate the effectiveness of our approach, we conducted two experiments. First, we assess the relevance of concepts that changed against those that remained unchanged between MeSH versions. Second, we compare the relevance of concepts between retracted and non-retracted papers in PubMed. We used the Mann-Whitney test to compare relevance changes; p-values below $0.05$ are considered statistically significant.
@@ 3 Results
This section presents results assessing the relevance of MeSH concepts using a multi-aspect model. It includes an analysis of the evolution of the MeSH terminology, highlighting key concepts at different hierarchy levels. Two use cases are explored: i) comparing the relevance of concepts that changed versus those that remained unchanged, and ii) comparing concepts used to annotate retracted papers versus those used for non-retracted papers.
@@ 3.1 MeSH concept relevance trends
We calculated the relevance of MeSH concepts from 2014 to 2021 using the PubMed 2022 snapshot annotated with the MeSH 2022 version. Relevance scores for the different aspects — disruptiveness, influence, informativeness, and usefulness — as well as for the overall concept relevance were computed monthly. Figure 2a depicts the cumulative trends, i.e., the sum of the first level of the MeSH hierarchy (A, …, Z). Disruptiveness, influence, and usefulness decrease over time ($p<0.001$). Park and colleagues also corroborate the decreasing trend of disruptiveness, although their analysis was based on scientific articles, in contrast to ours, which is based on MeSH concepts [36]. On the other hand, informativeness, which measures the degree of uncertainty, tends to increase over time ($p<0.001$). Therefore, the degree of uncertainty also increases as more recent knowledge is added to PubMed.
@@
Figure 2: Trend of concept relevance in MeSH at level 1 and correlation of descriptors. a) The sum of MeSH categories. b) Trend of MeSH categories. c) Ranking of concepts at level 1. d) Correlation of aspects. Column names are dis: disruptiveness (disruption), infl: influence (centrality), info: informativeness (entropy), use: usefulness (category utility), rel: concept relevance (RRF fusion).
@@
Figure 2b shows the trend of the concept relevance at the first level in MeSH. All of the categories show a decreasing trend ($p<0.001$). The most useful concepts (Figure 2b - bottom right) in the first level are D (Chemicals and Drugs), followed by N (Health Care) and C (Diseases), indicating they encode the basic-level categories of PubMed. In influence, all categories show a decreasing trend ($p<0.001$). Concept C is the most influential (Figure 2b - top right), indicating that ‘Diseases’ is a central concept in PubMed. It is followed by the concepts D and G (Phenomena and Processes), which include chemical, and physiological phenomena descriptors.
@@
Trends in disruptiveness are mostly decreasing ($p<0.001$), though categories M and H are not statistically significant. J shows an increasing trend ($p<0.001$). On the other hand, K shows a growing trend, while I shows no significant trend. Concept C is also the most disruptive in the first level of the MeSH hierarchy (Figure 2b - top left), being more than twice as disruptive as the second-ranked concept (D). Trends of informativeness are primarily increasing with a $p<0.001$. However, categories H, G, B, A, and D present a decreasing trend ($p<0.001$). Category C shows no trend (not statistically significant). Concept D has the highest informativeness score (Figure 2b - bottom left), followed by concept E (Analytical, Diagnostic and Therapeutic Techniques, and Equipment). However, its score gradually decreases, with a standard deviation of 0.003377. The trends of this aspect remain relatively constant over time, largely due to the narrow range of values across MeSH categories.
@@
Figure 2c shows the ranking of concepts in level 1. Most concepts retain their positions over time, except for Diseases (C), Health Care (N), Disciplines and Occupations (H), and I, which show some variation. Chemicals and Drugs (D) consistently ranks higher than other MeSH categories. The second position in this ranking alternates between ‘Diseases’ and ‘Health Care’, with the latter prevailing in five of eight years.